please help with polynomial.

x^3+125

thank you

Answer by nerdybill(7384)

Notice that:
x^3+125
.
Can be written as:
x^3+5^3
.
This is now a "sum of cubes" which is a special factor case:
http://www.purplemath.com/modules/specfact2.htm
a^3 + b^3 = (a + b)(a^2 – ab + b^2)
.
In our case,
a is 'x'
b is 5
.
So:
x^3+125 = (x + 5)(x^2 – 5x + 5^2)
x^3+125 = (x + 5)(x^2 – 5x + 25)